using the quadratic formula, what is the length and width of a rectangle if the length is 3 inches longer than twice the width and the area of the rectangle is 5 in^2?
Answer by oberobic(2304)

L = 2W + 3 (length = twice the width + 3)
A = L * W = 5 square inches
.
Substituting what we know.
A = (2W +3) * W = 5
2W^2 + 3W = 5
.
Subtract 5 from both sides
.
2W^2 + 3W - 5 = 0
.
We do not need to use the quadratic to solve this because it can be factored.
.
(2W + 5)(W - 1) = 0
.
So we have two candidate values for W: W = 1 and W = -5/2.
.
Since it is not possible to have a negative value for the width, W =1 is the candidate answer.
.
L = 2W + 3
L = 2(1) + 3
L = 5
.
Our candidate answer is:
L = 5
W = 1
.
Checking the work.
Is L*W = 5? yes.
So, we accept that answer.